OpenStudy (anonymous):

3! = A. 24 B. 9 C. 6 D. None of the above

OpenStudy (anonymous):

3! = 3 x 2 x 1 = 6

OpenStudy (gw2011):

3! = 3 x 2 x 1 = 6

OpenStudy (anonymous):

A factorial is the product of every integer starting at the given number to 1 (decrementing by 1). So 6! = 6 * 5 * 4 * 3 * 2 * 1, 24! = 24 * 23 * 22 ... and so on. So 3! = 3 * 2 * 1 = 6, as has been already given.